# ElementaryNet: A Non-Strategic Neural Network for Predicting Human Behavior in Normal-Form Games

### v2

@@ 6 Conclusions and Discussion
We see several promising directions for future work; we describe two here. First, in our experiments, we varied the strategic model and the potential functions, two pieces of the model with interpretable parameters. However, the response functions are still black-box neural networks. Developing methods for analyzing these responses would help to understand how the model is mapping from potentials to level-0 behavior.
Second, Zhu et al. (2024) presented an alternative method for improving existing behavioral game theory models, using a neural network to control the parameters of a quantal response model. Fusing their strategic models with our non-strategic models could be a productive way to make interpretable, yet highly predictive, models of human behavior.
@@ Appendix A Level-0 Heuristics
In this section, we enumerate the level-0 heuristic strategies used by Wright and Leyton-Brown (2019).
@@
Appendix B ElementaryNet Cannot Approximate Strategic Behavior
@@
$=\frac{\exp(\lambda\cdot u_{i}(a_{i}^{+},s_{-i}))}{\exp(\lambda\cdot u_{i}(a_{i}^{+},s_{-i})+\sum_{a_{i}\neq a_{i}^{+}}\exp(\lambda\cdot u_{i}(a_{i},s_{-i}))}$
@@
In their proof of Theorem 2.8, Wright and Leyton-Brown (2022) proved a stronger result that any convex combination of elementary models can be sensitive to their inputs in one of these two ways, but not both. {lemma} Let $f_{i}(G)=\sum_{k=1}^{K}w_{k}g_{i}^{k}(G)$ be a convex combination of elementary behavioral models. Then, either $f_{i}$ is not strongly dominance-responsive in the limit, it is not other responsive.
@@ D $x$ $x$
Appendix C Experiment Details
@@
We aggregated data from 12 experimental studies of human behavior in unrepeated normal-form games. The first ten were used in past work performing comprehensive evaluations of behavioral game theory models (Wright and Leyton-Brown 2017, 2019). The remaining two studies, by Fudenberg and Liang (2019) and Chui, Hartline, and Wright (2023), ran large-scale experiments on Amazon Mechanical Turk. Table˜1 lists details of these datasets.
@@
SW94 (Stahl and Wilson 1994) 10 400
SW95 (Stahl and Wilson 1995) 12 576
CGCB98 (Costa-Gomes, Crawford, and Broseta 2001) 18 1296
GH01 (Goeree and Holt 2001) 10 500
CVH03 (Cooper and Van Huyck 2003) 8 2992
HSW01 (Haruvy, Stahl, and Wilson 2001) 15 869
HS07 (Haruvy and Stahl 2007) 20 2940
SH08 (Stahl and Haruvy 2008) 18 1288
RPC08 (Rogers, Palfrey, and Camerer 2009) 17 1210
CGW08 (Costa-Gomes and Weizsäcker 2008) 14 1792
FL19 (Fudenberg and Liang 2019) 200 8250
CHW23 (Chui, Hartline, and Wright 2023) 24 4440
@@
To evaluate a single model on a single split of the data into a training, validation, and test set, we trained the model with a number of different hyperparameter settings, varying the initial settings of the QCH models’ parameters, the amount of L1 regularization applied to the neural networks’ weights, and the dropout probability applied to the neural networks. These hyperparameters are detailed in Table˜2. For each training run, we ran Adam (Kingma and Ba 2017) for 50,000 epochs with a learning rate of $3\cdot 10^{-4}$, using the projected gradient algorithm to impose simplex constraints where necessary.

### v3

@@ 6 Conclusions and Discussion
We see several promising directions for future work; we describe three here. First, we focused on interpreting the strategic model and the potential functions, two pieces of the model with interpretable parameters, but not the response functions, which are still black boxes. Future work could analyze these response functions by studying how they behave in simple cases, or by testing whether they continue to work well with additional restrictions.
Second, Zhu2024 presented an alternative method for improving behavioral game theory models, using a neural network to control the parameters of a quantal response model. Fusing their strategic models with ElementaryNet’s level-0 predictions could produce interpretable, yet highly predictive, models of human behavior.
Third, we focused on unrepeated, two-player normal-form games. These relatively simple games already evoke strategic behavior, but they do not include other elements such as sequential interactions, random chance, and imperfect information that are common in real-world settings. Extending our models to handle these elements is non-trivial, as it will require developing compelling definitions of non-strategic behavior beyond normal-form games and significantly modifying our architecture. Nonetheless, we are hopeful that the core idea of combining a rich model of non-strategic behavior with a structured model of strategic reasoning will be successful in these more complex settings.
Acknowledgments
This work was funded by an NSERC Discovery Grant, a CIFAR Canada AI Research Chair (Alberta Machine Intelligence Institute), and computational resources provided both by UBC Advanced Research Computing and a Digital Research Alliance of Canada RAC Allocation. Additionally, Greg d’Eon and Hala Murad were supported in part by funding from the UBC Advanced Machine Learning – Training Network, and Hala Murad was supported in part by funding from a UBC Work Learn International Undergraduate Research Award.
[…]
In these appendices, we include the following additional details:
  • Appendix A: formal definitions of additional level-0 heuristic strategies.
  • Appendix B: extending Theorem˜2 to handle negative utilities.
  • Appendix C: proving that ElementaryNet cannot represent broad categories of strategic reasoning.
  • Appendix D: additional details about experiments.
  • Appendix E: results in terms of absolute model performance.
@@ Appendix A Level-0 Heuristics
In this section, we enumerate the level-0 heuristic strategies used by JAIR.
@@
Appendix B Additional Proofs: GameNet with Negative Utilities
In Section 3, we argued that it is possible to extend the proof of Theorem 3.1 to handle negative utilities. We formalize this claim here. First, we restate the theorem, allowing for negative utilities.
Theorem 4.
Let $q_{1}(G)=QBR_{1}(\text{maxmax}_{2}(G);1,G)$. Let $\mathcal{G^{\prime}}$ be the set of games where all utilities are between $C_{max}$ and $C_{max}$ and all utilities differ by at least $C_{gap}$, where $C_{max},C_{gap}>0$ are arbitrary constants. Then, there exists an instantiation of GameNet’s feature layers that coincides with $q_{1}(G)$ for all $G\in\mathcal{G^{\prime}}$.
The proof is similar to the proof of Theorem 3.1.
Proof.
Like in the original proof, we first give a series of computations that produce $q_{1}(G)$ for all $G\in\mathcal{G^{\prime}}$. Let
$M_{c}^{\prime}$ $=\operatorname*{colmax}(U^{2}+C_{max}),$
$M_{*}^{\prime}$ $=\operatorname*{rowmax}(M_{c}^{\prime}),$
$B^{\prime}$ $=\operatorname*{relu}(M_{c}^{\prime}/C_{gap}-M_{*}^{\prime}/C_{gap}+1),$
$E^{\prime}$ $=\operatorname*{relu}(U^{1}+2C_{max}B^{\prime}-C_{max}),\text{and}$
$Q^{\prime}$ $=\textstyle\operatorname*{softmax}([\sum_{j}E_{i,j}^{\prime}]_{i=1}^{n}).$
Similar to before, $M_{c}^{\prime}$ is a matrix that contains the maximum utility that player 2 could realize by playing that action, plus $C_{max}$. $M_{*}^{\prime}$ is a constant matrix containing player 2’s maxmax value plus $C_{max}$. Because adding $C_{max}$ to all utilities does not change the maxmax action, $B^{\prime}$ is a matrix containing a column of ones for player 2’s maxmax action and zeros in all other columns. Then, consider the term $U^{1}+2C_{max}B^{\prime}-C_{max}$. In player 2’s maxmax column, this matrix contains player 1’s utilities plus $C_{max}$; because their utilities are greater than $C_{max}$, these entries are all positive. In the other columns, this matrix contains player 1’s utilities minus $C_{max}$; because their utilities are less than $C_{max}$, these entries are all negative. Thus, $E^{\prime}$ contains player 1’s utilities plus $C_{max}$ in player 2’s maxmax column, and zeros elsewhere. Finally, because constant offsets do not affect softmax operations, $Q^{\prime}$ is a vector containing player 1’s quantal best response to the maxmax action. Therefore, $Q^{\prime}=q_{i}(G)$.
Then, consider a GameNet model as before with three hidden layers, with two hidden units in the first two layers and one hidden unit in the final layer. In the first layer, setting $w^{1}_{1,2}=1$ and $b^{1}_{1}=C_{max}$ gives $H^{1,1}=U^{1}+C_{max}$, and setting $w^{1}_{2,6}=1$ and $b^{1}_{2}=C_{max}$ gives $H^{1,2}=M_{c}^{\prime}+C_{max}$. In the second layer, setting $w^{2}_{1,1}=1$ gives $H^{2,1}=U^{1}+C_{max}$, and setting $w^{2}_{2,2}=1/C_{gap}$, $w^{2}_{2,4}=-1/C_{gap}$, and $b^{2}_{2}=1$ gives $H^{2,2}=B^{\prime}$. In the last layer, setting $w^{3}_{1,1}=1$, $w^{3}_{1,2}=2C_{max}$, and $b^{3}_{1}=-2C_{max}$ gives $H^{3,1}=E^{\prime}$. The softmax operations at the end of the feature layers complete the proof. ∎
Appendix C Additional Proofs: ElementaryNet Cannot Approximate Strategic Behavior
@@
$=\frac{\exp(\lambda\cdot u_{i}(a_{i}^{+},s_{-i}))}{\exp(\lambda\cdot u_{i}(a_{i}^{+},s_{-i}))+\sum_{a_{i}\neq a_{i}^{+}}\exp(\lambda\cdot u_{i}(a_{i},s_{-i}))}$
@@
In their proof of Theorem 2.8, WLBnonstrategic proved a stronger result that any convex combination of elementary models can be sensitive to their inputs in one of these two ways, but not both. {lemma} Let $f_{i}(G)=\sum_{k=1}^{K}w_{k}g_{i}^{k}(G)$ be a convex combination of elementary behavioral models. Then, either $f_{i}$ is not strongly dominance-responsive in the limit, or it is not other responsive.
@@ D $x$ $x$
Appendix D Experiment Details
@@
We aggregated data from twelve experimental studies of human behavior in unrepeated normal-form games. The first ten were used in past work performing comprehensive evaluations of behavioral game theory models (WLB2017; JAIR). The remaining two studies, by FL19 and CHW23, ran large-scale experiments on Amazon Mechanical Turk. Table˜1 lists details of these datasets.
@@
SW94 (SW94) 10 400
SW95 (SW95) 12 576
CGCB98 (CostaGomes2001) 18 1296
GH01 (GH01) 10 500
CVH03 (CVH03) 8 2992
HSW01 (HSW01) 15 869
HS07 (HS07) 20 2940
SH08 (SH08) 18 1288
RPC08 (RPC08) 17 1210
CGW08 (CGW08) 14 1792
FL19 (FL19) 200 8250
CHW23 (CHW23) 24 4440
@@
Normalizing potential functions.
We also had some difficulty optimizing the potential functions in the ElementaryNet models with learned potentials. The problem was that when the potential function’s coefficients $\theta_{x}$ and $\theta_{y}$ were too large or too small, the resulting potentials had a variance much larger or smaller than 1. This caused the response functions to have inconsistent gradients and ultimately led to poor performance. To solve this problem, we divided the coefficients by $\sqrt{\theta_{x}^{2}+\theta_{y}^{2}}$ before applying the potential function, ensuring that the output potentials have a consistent scale. We found that this normalization step significantly improved performance.
Objective function.
Following the recommendation of dEon2024, we optimized the squared L2 error of the model’s predictions. Concretely, let $G=(A,u)$ be a game, let $\bar{p}(G)\in\Delta(A_{1})$ be the empirical distribution of actions that subjects played when acting as player 1, and let $f(G)$ be the model’s predicted distribution. Then, the squared L2 loss on this game is
$L(f(G),\bar{p}(G))$ $=\|f(G)-\bar{p}(G)\|_{2}^{2}$
$=\sum_{a\in A_{1}}(f(G)_{a}-\bar{p}(G)_{a})^{2}.$
We aggregated losses across multiple games by taking a weighted average, where each game was weighted in proportion to the number of observations it had in the dataset.
@@
To evaluate a single model on a single split of the data into a training, validation, and test set, we trained the model with a number of different hyperparameter settings, varying the initial settings of the QCH models’ parameters, the amount of L1 regularization applied to the neural networks’ weights, and the dropout probability applied to the neural networks. These hyperparameters are detailed in Table˜2. For each training run, we ran Adam (Adam) for 50,000 epochs with a learning rate of $3\cdot 10^{-4}$, using the projected gradient algorithm to impose simplex constraints where necessary.
@@
Appendix E Additional Results
In Section 5, rather than directly reporting the loss of each model, we reported differences between models’ losses. This increased statistical power by reducing variance caused by differences in the data split used to evaluate each replication of each model. However, this statistical power came at the expense of making it harder to gauge each model’s performance in absolute terms. In this section, we provide analogous figures showing absolute test losses:
  • Figure˜8 compares GameNet and ElementaryNet models with a variety of neural net widths and depths, and a QCHp strategic model.
  • Figure˜9 compares ElementaryNet models with 1 to 3 hidden layers and no strategic model.
  • Figure˜10 compares ElementaryNet models with alternative strategic models (QCH1, QCH2, and QCH3).
  • Figure˜11 compares ElementaryNet models with alternative potential functions: one with a single potential fixed to the “own” potential function, and one with all four of the fixed potential functions that we defined.
Figure 8: ElementaryNet and GameNet level-0 models trained with a QCH-Poisson strategic model.
Figure 9: ElementaryNet models with no QCH model.
Figure 10: ElementaryNet level-0 models with various QCH models.
Figure 11: ElementaryNet level-0 models with various potential functions.